Calculate caloricity of the 150 g of milk, which contains 25% of protein, 2,5% of fats and 23% of carbohydrates.


1
Expert's answer
2021-11-23T01:21:02-0500

150×0.25=37.5g

150×0.025=3.75g

150×0.23=34.5g